OpenStudy (anonymous):

Just wondering if anyone could explain how/why the answer to this is still -17. Thanks! Suppose v is in the interval (π/2 , π), with tan v = −17. Find the exact expression for the indicated quantity. tan(v + 3π)

OpenStudy (loser66):

\[tan (a +b) =\frac{tan a+tan b}{1-tana tanb}\]your a =v your b =3\(\pi\), and tan 3\(\pi\)=0, therefore, yours = \[\frac{tan v}{1} =tan v = -17\]

OpenStudy (anonymous):

Thanks!